OpenStudy (anonymous):

Can someone please help me with this question A number is increased by forty percent, or three hundred sixty. What is the original number?

OpenStudy (anonymous):

Original number = x X+360 = y 1.4 * x = y 0.4x = 360 X = 360/0.4

OpenStudy (anonymous):

Maybe it's more clear if I write it like this: X+360 = x+0.4x 360 = 0.4x 360/0.4 = x

OpenStudy (anonymous):

I really don't understand so want to know and answer to :)